4 statistics professors and 6 chemistry professors are
available to be advisors to a student organization. The
student organization needs three of the professors to be
advisors. If each professor has an equal chance of being
selected, what is the probability that both professors are
chemistry professors?

Answer by stanbon(75887)

4 statistics professors and 6 chemistry professors are
available to be advisors to a student organization. The
student organization needs three of the professors to be
advisors. If each professor has an equal chance of being
selected, what is the probability that both professors are
chemistry professors?
-------
# of ways to succeed: 6C2*4C1 = 60 ways
-----
# of possible outcomes:: 10C3 = 120 ways
--------------
P(2 chem profs and one stat prof) = 60/120 = 1/2
==============
Cheers,
Stan H.
=================